# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ 2.2.1 First Portfolio Construction
where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for $j=1,...,p$. The problem (2.1) then becomes
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$
This is a constrained convex optimization problem. By Lagrange multiplier method and the definition of $\bm{w}^{*}$ and $\mathcal{A}_{t}$, we have the explicit solution $\bm{w}_{1}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{1}^{-1}\left(\gamma\bm{\mu}_{1}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\bm{\mu}_{1}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\mathbf{1}}\mathbf{1}\right)$, and $\bm{w}^{*}_{\mathcal{A}_{1}^{c}}=\mathbf{0}$, by sparsity assumption, where $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, $\mathbb{B}=\text{diag}(\beta_{1},...,\beta_{p})$.
As discussed earlier, we use the LLA algorithm for SCAD, and $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ is the initial and Lasso solution. In the following lemma, we show the consistency of the initial Lasso weight estimator.
@@
For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$, where $C_{2}$ is defined in Assumption (A2). Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}=1}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{Lasso}\|\bm{w}\|_{1}$ (2.7)
with $\lambda_{Lasso}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right),\quad\quad\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$. In other words, if the portfolio has sufficient sparsity that $s_{1}=o(\sqrt{\frac{n}{\log p}})$, the estimation error can be ignored. Our error rate also matches the rate of Lasso estimator for linear regression. For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+(\bm{\beta}_{{\mathcal{A}_{1}}}\odot\bm{w}_{{\mathcal{A}_{1}}})^{\top}\bm{w}_{\mathcal{A}_{1}}.$ (2.8)
The following theorem states that we can find the oracle estimator after two iterations by implementing the LLA algorithm, and it also provides the convergence rate in $\ell_{\infty}$-norm.
@@
Suppose that Assumptions (A1)-(A4) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), we choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 1 states that under mild conditions, our estimator agrees with the oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle},\beta}$ with high probability when $k$ diverges to infinity. Hence, we only study the property of non-convex penalty in the following sections.
@@ 2.2.2 Portfolio Reallocation
At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$. And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$. The form of transaction costs $C_{t}(\bm{w})$ is thus given by
$\mathbf{C}_{t}(\bm{w})=[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1})$
@@
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t}+[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1}),\quad\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$ (2.9)
With $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, the optimization problem has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\quad\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.10)
Here, we need to modify Assumption (A4) and propose the following revised assumption:
  • (A4’)
    Suppose the theoretical minimizer $\bm{\delta}_{t}^{*}$ is sparse with cardinality $s_{t}=\|\bm{\delta}^{*}_{t}\|_{0}$ and the minimal signal strength of $\bm{\delta}_{t}^{*}$ satisfies that $\|\bm{\bm{\delta}}_{\mathcal{A}_{t}}^{*}\|_{\min}>(a+1)\lambda$;
Similarly, using the Lagrange multiplier method, we have that $\bm{\delta}_{t}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{t}^{-1}(\gamma\tilde{\bm{\mu}}_{t}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\mathbf{1}}\mathbf{1})$ and $\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}$, where $\tilde{\bm{\mu}}_{t}=\bm{\mu}_{t}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$.
@@
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,$ (2.11)
where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the t period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$. By incorporating the transaction costs in the mean variance problem and regularizing the eventual rebalancing events, our portfolio achieves better Sharpe ratio and less transaction costs, which is shown in the numerical studies.
@@
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\tilde{\bm{\Sigma}}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}.$ (2.12)
@@
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we select $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso }}.$
@@
    If $\|\bm{w}^{+}_{t-1}\|_{1}<\infty$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
@@
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq\infty$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Theorem 2 implies that in portfolio reallocation, the LLA algorithm initialized at a Lasso solution yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error. Theorem 2 highlights the impact of different estimation methods employed in the previous $t-1$ stages on the performance of our proposed method in stage $t$. Specifically, if regularized estimation methods, such as the proposed CAPE or Lasso, were consistently applied in the previous $t-1$ stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity. This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage. Conversely, if the estimation methods used in the earlier stages fail to produce sparse estimates, leading to a non-sparse $\bm{w}^{+}_{t-1}$, our method can still ensure that $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\bm{\delta}^{*}_{t}$ with probability approaching 1. However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$. Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE method is consistently applied across the previous $t-1$ stages.
@@
Define $\max\{s_{1},\ldots,s_{t}\}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE to make estimation in the past $t-1$ stage. Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
@@
2.3 CAPE with Proportional Transaction Cost

### v2

@@ 2.3.1 First Portfolio Construction
where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for j=1,,pj=1,...,p. The problem (2.4) then becomes
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0}.$ (2.11)
This is a constrained convex optimization problem. By applying the method of Lagrange multipliers and using the definitions of $\bm{w}^{*}_{1}$ and the active set $\mathcal{A}_{1}$, we obtain the explicit solution
$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{w}_{\mathcal{A}^{c}_{1}}^{*}=\mathbf{0}.$
Here, $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$. The explicit solution above is derived using the method of Lagrange multipliers and the sparsity structure of the optimizer. For a detailed derivation, we refer the reader to the proof of Theorem 1 in the supplementary material.
As discussed earlier, we employ the LLA algorithm to solve CAPE-S and use the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ as the initial estimate required by the LLA algorithm. The following lemma establishes the consistency of the CAPE-L.
@@
For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1}$ (2.12)
with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
In Lemma 1, the estimator $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ defined in Equation (2.12) essentially applies the $\ell_{1}$ penalty to the penalty term introduced in Equation (2.6). Following the naming convention used in Equation (2.8), this class of estimators can be referred to as the Cost-Aware Portfolio Estimator under Lasso Penalty (CAPE-L) at the first portfolio construction stage. Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$. In other words, if the portfolio has sufficient sparsity that $s_{1}=o(\sqrt{\frac{n}{\log p}})$, the estimation error can be ignored. Our error rate also matches the rate of Lasso estimator for linear regression.
For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}.$ (2.13)
The following theorem states that the oracle estimator can be obtained after two iterations by implementing the LLA algorithm to solve CAPE-S, and it also provides the convergence rate in the $\ell_{\infty}$-norm.
@@
Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some sufficiently large constant $M$. Let $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ denote the solution obtained from the LLA algorithm 1 with the SACD penalty, initialized at $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{Lasso},\beta}$. If $s_{1}=O\!\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{w}^{*}_{1}\|_{\min}>(a+1)\lambda_{1}$, then $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{oracle},\beta}$ within two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[s_{1},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\mathrm{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\!\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 1 establishes that, under mild regularity conditions, the proposed CAPE-S estimator attains the same performance as the oracle estimator $\widehat{\bm{w}}_{1}^{\mathrm{oracle},\beta}$ after only two iterations of the LLA algorithm. In particular, when the portfolio dimension $k$ diverges, the two estimators coincide with high probability, ensuring that CAPE-S inherits the oracle’s optimality properties in large-scale settings. This oracle property provides the theoretical foundation for subsequent results on risk and return estimation.
@@ 2.3.2 Portfolio Reallocation
At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$. And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$. The form of transaction costs $C_{t}(\bm{w})$ is thus given by
[…]
@@
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|\leq s_{0}.$ (2.14)
Similarly, using the Lagrange multiplier method, we have that
$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}.$
@@
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,$
where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the $t$ period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$. By incorporating the transaction costs in the mean variance problem and regularizing rebalancing events, our portfolio achieves a higher post-cost Sharpe ratio and reduced transaction costs, as demonstrated in numerical studies.
@@
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}.$ (2.15)
where $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$.
Lemma 2.
Suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$ with $C_{1}$ defined in Assumption (A1), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.16)
with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
@@
Suppose that Assumptions (A1)-(A3) hold, and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1). Let $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, and assume that $\left\|\bm{\delta}^{*}_{t}\right\|_{\min}>(a+1)\lambda_{t}$. Choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some sufficiently large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$, as defined in (2.16). Then, we have the following results:
@@
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
@@
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 2 and Lemma 2 have implied that in portfolio reallocation, the LLA algorithm initialized at CAPE-L yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error. Theorem 2 highlights the impact of different estimation methods employed in the previous t1t-1 stages on the performance of our proposed method in stage $t$. Specifically, if regularized estimation methods, such as the proposed CAPE-S in (2.8) or CAPE-L in (2.12), were consistently applied in the previous t1t-1 stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity. This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage. Conversely, if the estimation methods used in the earlier stages fail to produce sparse estimates, leading to a non-sparse $\bm{w}^{+}_{t-1}$, our method can still ensure that $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\bm{\delta}^{*}_{t}$ with probability approaching 1. However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$. Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE-S method is consistently applied across the previous t1t-1 stages.
@@
Define $s_{0}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE-S to make estimation in the past t1t-1 stage. Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
@@
2.4 CAPE-S with Proportional Transaction Cost